# PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation

### v1

@@ 6.5 Error Analysis
To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize three types of error in Table 4). For each example, we report the original PLS sentences, the corresponding scientific abstract with retrieved knowledge (from medical Textbooks and StatPearls), model-generated questions based on the extracted answers, and the QA model’s responses.
In the first case, the extracted answer is “pregnancy,” but the QA model returns “cycle control” from the abstract due to the absence of this term in the retrieved content. Since this sentence is labeled as “explanation” which requires additional knowledge for verification, the error suggests potential misclassification between simplification and explanation. The second example shows how noisy retrieved information can affect evaluation accuracy, as the QA model’s answer is irrelevant to the extracted answer. In the final case, an unanswerable question generated by the QG model (e.g., regarding “Government”) reveals limitations in QA-based evaluation, as the question remains unanswered even with retrieval. Additionally, we acknowledge that some PLS summaries or sentences may not generate any questions, resulting in empty question sets. Addressing these challenges requires further improvements in classification, retrieval, and question generation.
@@ 7 Discussion and Conclusion
Automated PLS generation has great potential for improving health information accessibility, but factuality evaluation remains challenging due to elaborative explanations. To address this, we introduce PlainQAFact, the first PLS factuality evaluation framework. PlainQAFact classifies sentence types, retrieves domain-specific knowledge, and utilizes a QA-based approach to assess the factual consistency of both source-simplified and elaborative sentences. By leveraging open-source models, PlainQAFact ensures both accessibility and computational efficiency.
To support granular factuality evaluation, we construct PlainFact, the first expert-annotated dataset for PLS, covering factuality type, functional role, and sentence alignment. Using PlainFact, our experiments demonstrate that PlainQAFact outperforms existing factuality metrics, particularly in handling external elaborative content. However, applying sentence-level evaluation to large datasets can be computationally expensive. For extensive data, we recommend using summary-level evaluation as a more efficient alternative while maintaining comparable performance.
Although our study focuses on biomedical PLS generation, the retrieval framework is adaptable to other domains by substituting domain-specific knowledge bases. Our findings highlight the effectiveness of retrieval-augmented QA for factuality evaluation, suggesting its broader applicability beyond PLS tasks. Future research should explore its integration into diverse summarization applications, particularly in LLM-generated text, where source coverage is often incomplete.
By introducing a fine-grained and explainable evaluation framework, our work establishes a foundation for advancing QA-based factuality metrics. We hope PlainQAFact will contribute to the development of more reliable and interpretable factuality assessment methods in NLP.
Limitations
We acknowledge that using an LLM as an answer extractor may yield non-deterministic outputs, even when setting the temperature to 0.01 for the Llama 3.1 8B Instruct model. For example, in our retrieval effectiveness evaluation (Table 3), we run PlainQAFact four times to examine the impact of LLM-based answer extraction on the final factuality prediction. The resulting scores were 50.60, 50.64, 50.57, and 50.63, with a standard deviation of 0.027. We consider this variability to be minimal, demonstrating that our factuality evaluation metric shows robust performance in the plain language summarization evaluation task.
Our study’s major limitations are as follows: 1) Although we provide substitute solutions for classifying input summaries or sentences, our pre-trained classifier is limited to the biomedical domain and specifically designed for plain language summarization tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts may focus on developing domain- and task-specific classifiers You et al. (2024a) for factuality evaluation. 2) In the knowledge retrieval phase, we currently incorporate only Textbooks and StatPearls as external resources. This limited selection may not be sufficient for evaluating the factuality of plain language summaries. Further investigation into additional relevant knowledge bases, such as Wikipedia and PubMed, could enhance performance in explanation-needed factuality evaluation tasks. 3) For retrieval-augmented QA, we directly concatenate the scientific abstract with the retrieved information to form the reference for QA. However, we do not apply knowledge filtering to remove irrelevant information from external resources, which may introduce noise into the subsequent QA process for assessing answer-level factual correctness.
@@ Appendix A Dataset Annotation Protocol
Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators. Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the PLS and its corresponding scientific abstract. For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
@@
Each annotator will annotate 40 PLS-abstract pairs to ensure each sentence of the PLS has two sets of annotations from different people. For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
$<$TO BE ANNOTATED$>$ External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
$<$TO BE ANNOTATED$>$ Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
$<$TO BE ANNOTATED$>$ Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5). For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3. Use commas to separate multiple indices.
@@
Corresponding Abstract $<$s10_1$>$Acute cardiovascular events represent a therapeutic challenge. $<$s10_2$>$Blood pressure lowering drugs are commonly used and recommended in the early phase of these settings.
@@
Corresponding Abstract $<$s15_16$>$For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence). $<$s15_17$>$Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence). $<$s15_20$>$For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
Table 5: Examples of our curated dataset. Need External Information feature represents whether a plain language sentence is a simplification or an explanation. A label of “yes” indicates that the sentence is an explanation and requires additional elaborative information beyond the source abstract to verify its factuality. Conversely, a label of “no” shows that the sentence is a simplification that can be validated using only the source abstract.
@@ Appendix B Dataset Examples
In Table 5, we presents three representative examples from PlainFact. Each example is annotated with five features: a plain language sentence, an indicator of whether the sentence is simplification or explanation, its category, its relation, and the corresponding abstract. All plain language sentences are factual, as they were written by the authors from the Cochrane database. “Need External Information?” feature specifies whether a sentence can be validated solely by the abstract. A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa. The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the PLS; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure. Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
@@ Appendix C LLM Prompts
We utilize two types of prompts to guide LLMs through two stages of factuality evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (Appendix A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factuality evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval Fabbri et al. (2022a), we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factuality of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 8B Instruct as a factuality judge.
@@ Appendix D Detailed Experiment Settings
All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory. We employ the Natural Language Toolkit (NLTK) [^6] https://www.nltk.org/ for PLS tokenization and sentence segmentation, utilizing the punkt package. For classifier fine-tuning, we set the random seed to 42 to split the dataset into training, validation, and test sets. We tune the PubMedBERT-base model for three epochs with early stopping under the validation set. The same seed (42) is also used to sample 200 PLS-abstract pairs from the CELLS test set for the comparison in Figure 2. The QA model used in our PlainQAFact is downloaded from QAFactEval Fabbri et al. (2022a); however, we set its maximum input length to 512 (from 364) tokens to incorporate more source context. Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens.
Component Method Choice Benchmark
PlainQAFact 55.8 (0.12)
Classifier Learned classifier -
GPT-4o 59.1 (0.05)
No (retrieve for all) 60.1* (0.18)
Answer Extraction Llama3.1 8B Instruct -
GPT-4o 54.8 (0.15)
PromptRank 48.4*
NP Chunks 50.5*
Retrieval Source Abs + TB + SP -
TB + SP 47.9* (0.11)
Abs (no retrieval) 55.2 (0.07)
Answer Overlap BERTScore -
LERC 32.9* (0.18)
Table 6: Ablation study of PlainQAFact on PlainFact in the summary level, analyzing the impact of individual components in PlainQAFact. Abs: abstract; TB: Textbooks; SP: StatPearls. Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
Appendix E Summary-Level Ablation Study of PlainQAFact
Similar to the sentence-level ablation studies in Table 2, we conduct experiments to evaluate the impact of various components in PlainQAFact at the summary level using PLSs as inputs. Although the overall evaluation score using GPT-4o as the classifier is 3.3% higher than that of PlainQAFact, the paired t-test (p<0.05) does not provide sufficient statistical evidence to conclude that GPT-4o significantly outperforms our learned classifier. This finding indicates that GPT-4o’s performance is comparable with our learned classifier, consistent with the sentence-level results.
Notably, in the summary-level evaluation, using only scientific abstracts as references for QA-based evaluation yields performance that is competitive with using retrieved external knowledge. We assume that summary-level retrieval may not consistently produce highly relevant knowledge snippets and might even introduce noise that undermines factuality evaluation accuracy.
Additionally, omitting the classification stage and retrieving information for all summaries produces the best performance among all settings. Although this approach increases the runtime to approximately 30 minutes compared to using the learned classifier for 200 summaries, the performance improvement is significant. Therefore, we recommend either employing more targeted knowledge bases for summary-level retrieval, compressing abstracts to include more relevant information, or retrieving for all input summaries when conducting summary-level factuality evaluation with PlainQAFact.
Figure 5: Evaluation performance across metrics for explanation sentences under PlainFact. We select only those summaries containing at least five explanation sentences, resulting in a total of 116 summaries. Note that this experiment is conducted at the summary level. A higher change percentage indicates a metric is more sensitive to the explanation information, implying that such metrics may not effectively verify the factuality of PLS. Ideally, we expect the score change percentage to increase slowly. Among the metrics evaluated, only AlignScore, SummaC, and our method show this expected pattern.
Appendix F Explanation Sentence Removal
To further assess the performance of existing factuality metrics on elaborative explanation sentences in PLSs, we conduct experiments in which these explanation sentences were iteratively removed from the summaries. Figure 5 illustrates the percentage change in scores relative to the baselines as more explanation sentences are removed. Notably, PlainQAFact can even predict 100% factual ratings for the given PLS when more explanation information is removed, whereas other metrics struggle to do so. Although our framework shows moderate performance in Figure 5, with a gradual increase in score as explanation sentences are removed—we believe that our retrieval-augmented, QA-based factuality evaluation framework more effectively assesses the factuality of plain language sentences that require elaborative explanations for verification.

### v2

@@ 5.4 Error Analysis
To analyze cases where PlainQAFact fails on the PlainFact benchmark, we categorize one correct example and three types of errors in Table 5. For each case, we present the original plain language sentence, the corresponding scientific abstract with retrieved knowledge (from medical textbooks and StatPearls), the model-generated questions based on extracted answers, and the QA model’s final responses. Each question is generated based on the extracted answer and its corresponding plain language sentence.
In the correct case, the QG model generates a question from the plain language sentence and the extracted answer “SSB.” The QA model then provides the correct answer using the retrieved knowledge rather than the original abstract, demonstrating PlainQAFact ’s effectiveness in evaluating explanation sentences.
In the first failure case, the extracted answer is “pregnancy,” but the QA model returns “cycle control” from the abstract because the correct term is missing in the retrieved content. Since this sentence is classified as an “elaborative explanation,” which requires external knowledge for verification, this error points to a potential misclassification between simplification and explanation by the fine-tuned classifier. The second failure illustrates how noisy retrieved knowledge can impair evaluation. The QA model provides an irrelevant answer from the retrieved content that does not match the extracted gold answer. The third case shows an unanswerable question generated by the QG model (e.g., regarding “Government”), highlighting a limitation of QA-based factual consistency evaluation. Some questions remain unanswered even with retrieval. Additionally, we acknowledge that some plain language sentences may not generate any questions, resulting in empty question sets. Addressing these challenges requires further improvements in classification, retrieval, and domain-specific question generation.
@@ 6 Discussion and Conclusion
Our study advances the assessment of hallucinations in the field of PLS by supplying both the first domain expert-annotated biomedical PLS benchmark PlainFact and a novel retrieval-augmented QA-based factual consistency evaluation metric PlainQAFact. Unlike existing biomedical domain corpora that either lack fine-grained labels for elaborative content [54] or focus solely on text simplification (e.g., FactPICO[9]), PlainFact captures sentence-level distinctions, including simplification versus explanation, functional roles, and explicit alignment to source sentences. This granularity not only enables precise error analysis of hallucinations introduced for clarity, but also provides a reusable framework for researchers in other disciplines (e.g., legal, technical) to replicate our annotation protocol and build high-quality factual consistency datasets tailored to their specialized texts. By offering these data and detailed annotation protocol (A), we anticipate community-driven extensions, such as multilingual adaptations or integration with domain-specific ontologies that will democratize development of trustworthy summarization models.
Building on this resource, PlainQAFact differentiates from prior QA-based metrics (e.g., QAFactEval[21], QuestEval[50]) by first classifying whether a sentence requires external verification given elaborative explanations are often ignored by metrics relying solely on source text, and improve the evaluation efficiency without retrieving for every plain language summary instance. Then, retrieving targeted domain knowledge for each elaborative explanation before posing and answering fact-checking questions. The ablations demonstrate that selective retrieval yields substantial gains in Kendall’s $\tau$, Pearson correlation, and AUC-ROC over both end-to-end LLM judges[41] and alignment-based scorers[19], highlighting that judicious incorporation of external evidence is essential for robust factual consistency assessment. Beyond its empirical strengths, PlainQAFact serves as a plug-and-play evaluation metric for future summarization systems. Developers can leverage our classifier to classify sentences, apply retrieval only where necessary, thus containing computational costs and obtain interpretable question–answer pairs that manifest non-factual hallucinations. More importantly, PlainQAFact is developed with the open-source backbone model (i.e., Llama 3.1), ensuring broad use with transparency, reproducibility, and fewer computational costs. Looking forward, extending our proposed evaluation method to interactive human–AI workflows, where clinicians or subject-matter experts verify QA outputs on-the-fly, could further enhance trust and adoption in high-stakes domains, including medicine, healthcare, or law.
Despite these promising results, we consider the following limitations of this study. (1) Our error analysis reveals that misclassifications, particularly between source simplifications and elaborative explanations, can lead to retrieval mismatches and unanswerable QA prompts. These issues underscore the need for more nuanced classification and question generation modules. Although we provide substitute solutions for classifying input summaries or sentences, our fine-tuned classifier is limited to the biomedical domain and specifically designed for PLS tasks, fine-tuned through PlainFact. We acknowledge that the dataset used for model fine-tuning is insufficient, resulting in only moderate classification accuracy. Given the scarcity of high-quality human-annotated data for classifying source simplification and elaborative explanation information, future efforts are needed on developing domain-specific (e.g., law, healthcare, social science, etc.) classifiers [55] for factual consistency evaluation. (2) While our experiments indicate that our retrieval-augmented evaluation metric can improve factual consistency assessment in most of the settings, especially in elaborative explanation evaluation, the computational time increases compared to NLI-based evaluation metrics and LLM-based evaluators. The trade-off between evaluation precision and efficiency suggests that further optimization will be beneficial. We suggest to balance wisely based on the trade-offs we report in this study regarding the evaluation time and accuracy.
@@ Appendix A Dataset Annotation Protocol
Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators. Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the plain language summary and its corresponding scientific abstract. For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
@@
Each annotator will annotate 40 summary-abstract pairs to ensure each sentence of the plain language summary has two sets of annotations from different people. For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
<<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
<<TO BE ANNOTATED>> Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
<<TO BE ANNOTATED>> Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5). For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3. Use commas to separate multiple indices.
@@
Corresponding Abstract <<s10_1>>Acute cardiovascular events represent a therapeutic challenge. <<s10_2>>Blood pressure lowering drugs are commonly used and recommended in the early phase of these settings.
@@
Corresponding Abstract <<s15_16>>For treatment failure, the difference between doxycycline and tetracycline is uncertain (very low-certainty evidence). <<s15_17>>Doxycycline compared to tetracycline may make little or no difference in resolution of fever within 48 hours (risk ratio (RR) 1.14, 95% confidence interval (CI) 0.90 to 1.44, 55 participants; one trial; low-certainty evidence) and in time to defervescence (116 participants; one trial; low-certainty evidence). <<s15_20>>For most outcomes, including treatment failure, resolution of fever within 48 hours, time to defervescence, and serious adverse events, we are uncertain whether study results show a difference between doxycycline and macrolides (very low-certainty evidence).
Table 6: Examples of our curated dataset. Need External Information feature represents whether a plain language sentence is a simplification or an explanation. A label of “yes” indicates that the sentence is an explanation and requires additional elaborative information beyond the source abstract to verify its factual consistency. Conversely, a label of “no” shows that the sentence is a simplification that can be validated using only the source abstract.
@@ Appendix B Dataset Examples
In Table 6, we presents three representative examples from PlainFact. Each example is annotated with five features: a plain language sentence, an indicator of whether the sentence is simplification or explanation, its category, its relation, and the corresponding abstract. All plain language sentences are factual, as they were written by the authors from the Cochrane database. “Need External Information?” feature specifies whether a sentence can be validated solely by the abstract. A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa. The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure. Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
@@ Appendix C LLM Prompts
We utilize two types of prompts to guide LLMs through two stages of FC evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruc model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the FC evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the FC of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 8B Instruct as a FC judge.
For the LLM-based perturbation of PlainFact and CELLS datasets, we follow the protocol of APPLS [12] on faithfulness criteria.
System
You are a data transformation assistant. You will receive a sentence from a biomedical literature. You will generate a new version of the given sentence based on the following rules for faithfulness perturbations:
1. Number Swap
Locate any numeric value(s) in the sentence and swap them with different numeric value(s).
Example: "infected more than 59 million people" -> "infected more than 64 million people"
2. Entity Swap
Locate a key entity (e.g., virus name, drug name, organization) in the sentence and swap it with a different entity.
Example: "coronavirus 2 (SARS-CoV-2)" -> "canine adenovirus (CAV-2)"
3. Synonym Verb Swap
Identify a key verb in the sentence and replace it with a near-synonym or related verb that changes the nuance or meaning slightly.
Example: "killed more than one of them" -> "stamped out more than one of them"
4. Hypernym/Antonym Swap
Select a word and replace it with either a hypernym (a more general term) or an antonym (opposite meaning), as appropriate.
Example (antonym): "killed more than one of them" -> "saved more than one of them"
Example (hypernym): "dog" -> "animal" (if relevant)
5. Negation
Negate a key part of the sentence to flip its meaning.
Example: "has infected more than 59 million people" -> "hasnt infected more than 59 million people" or "has not infected more than 59 million people"
Your task:
Read each sentence, generate a new sentence based on one of the five types of perturbation stratigies (Number Swap, Entity Swap, Synonym Verb Swap, Hypernym/Antonym Swap, Negation) above.
Return the perturbation sentence.
Do not change the rest parts of the sentence except for the perturbation content. For example, the original sentence is "The skin patch and the vaginal (birth canal) ring are two methods of birth control." The perturbation sentence should be "The skin patch and the vaginal (birth canal) ring are five methods of birth control."
User
Sentence: <input>
Perturbation sentence: <sentence>
Example 5: Prompt of GPT-4o for faithfulness perturbation
@@ Appendix D Detailed Experiment Settings
All experiments we conduct are under one NVIDIA A100 GPU with 40 GB GPU memory. We employ the Natural Language Toolkit (NLTK) [^5] https://www.nltk.org/ to split the plain language summaries into sentences using the punkt package. For the classifier fine-tuning, we set the random seed to 42 to split PlainFact into training, validation, and test sets. We tune the PubMedBERT-base model for three epochs with early stopping under the validation set. The final fine-tuning accuracy of the classifier on the test set is 0.77. The same seed (42) is also used to sample 200 summary-abstract pairs from the CELLS test set for the comparison in Table 2.
The QA model used in our PlainQAFact is downloaded from QAFactEval [39]. However, we set its maximum input length to 512 (from 364) tokens to incorporate more source context. Similarly, to ensure that the answers extracted during the AE stage are valid (i.e., within the 512-token limit to maintain consistency with the subsequent QA model), we configure the LLMs’ input lengths to 512 tokens. In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (D).
We apply the default setting (i.e., RoBERTa-base) for AlignScore [19] and SummaC-Conv for SummaC [16]. For GPT-4o experiments as an factual consistency evaluator, we set the temperature as 1.0 in this paper.
Appendix E Statistical Testing
Following SummaC [16] and QAFactEval [39], we test whether our proposed metric achieves statistically significant improvements over other methods. Given Kendall’s $\tau$ and Pearson correlations do not show the discriminative feature of continuous scores in various evaluation metric, we run a systematic evaluation on four datasets (CELLS, FareBio, FactPICO, and PlainFact) through bootstrap resampling [56] on the AUC-ROC results. We compare our best metric to other methods using the confidence intervals as significance level of 0.05 and apply Bonferroni correction [57] following SummaC [16]. Our results show that PlainQAFact achieves statistically significant improvements over QuestEval [50], SummaC [16], and DAE [48] on the CELLS and PlainFact datasets. As shown in Figure 2, PlainQAFact also outperforms other automatic metrics in evaluating elaborative explanation sentences, except for QAFactEval [39]. This highlights the effectiveness of PlainQAFact for biomedical PLS tasks that include external explanations. On the other hand, when evaluating elaborative content in the FactPICO and FareBIO datasets (Figures 3 and 4), we do not find statistically significant differences across metrics. While some existing metrics perform similarly in certain settings, our metric consistently shows better performance across all three evaluation criteria.
Figure 3: Overall performance on summaries containing added information (i.e., elaborative explanations) from FactPICO [9]. The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.2, 0.2, and 3.3, respectively, based on five runs of each metric (details in  E).
Figure 4: Overall performance on summaries with elaborative explanations from FareBio [47]. The std. for PlainQAFact, Llama 3.1, and GPT-4o are 0.03, 0.3, and 5.7, respectively, computed over five runs (details in  E.)
Appendix F Summary-Level Explanation-Only Evaluation
Similar to our sentence-level explanation annotations in PlainFact, FactPICO [9] and FareBio [47] datasets also provide human-annotated explanation information in plain language summaries. According to the results shown in Section 5.2, we also assess the performance of five factual consistency evaluation metrics under FactPICO and FareBio datasets.
In the FactPICO, the elaborative explanation is defined as “added information.” We first select summaries that include added information, and then we label a summary as “non-factual” if it contains any non-factual added information as determined by annotators. We generate a summary-level explanation-only FactPICO dataset consisting of 231 summary-abstract pairs. As shown in Figure 3, AlignScore achieves the best performance across all three evaluation criteria. However, since the summaries in the FactPICO dataset are generated using LLMs, the factual consistency between the generated summaries and the original source abstracts is not guaranteed. For instance, we treat a summary as “factual” indicates that all the added information is factual, but it does not necessarily reflect the factual consistency of other content. Therefore, the results presented in Figure 3 may become biased.
For FareBio [47], we collect explanation sentences based on both “faithfulness” and “factual hallucination” labels in the original dataset. As shown in Figure 4, PlainQAFact achieves the best performance compared with all other metrics, showing a clear advantage over GPT-4o. Since the FareBio dataset provides sentence-level annotations of factual consistency for each sentence in the plain language summaries, the results are more reliable than those from FactPICO.